OpenStudy (khaakenajaf):

i understand the formulas for similar figures but i have a confusion, if x1=side of small triangle and x2=side of big triangle then how to choose x1/x2 or x2/x1 ?i dont understand which shape big or small comes in the numerator and demominator

OpenStudy (amistre64):

doesnt matter, as long as you keep the side placements consistent, you are looking for a ratio. 6:5 is the same as 5:6 when comparing similar parts

OpenStudy (anonymous):

bye @amistre64 :)

OpenStudy (amistre64):

ciao?

OpenStudy (anonymous):

ok see u later :)

OpenStudy (amistre64):



OpenStudy (amistre64):

\[\frac{a_1}{a_2}=\frac{b_1}{b_2}\]or \[\frac{a_1}{b_1}=\frac{a_2}{b_2}\]or \[\frac{b_2}{b_1}=\frac{a_2}{a_1}\]

OpenStudy (amistre64):

and i spose:\[\frac{b_2}{a_2}=\frac{b_1}{a_1}\] just to complete the permutations

OpenStudy (khaakenajaf):

@amistre64 thanks for your efforts i get the formulas and the drawing but still a thing i am not getting. in these shapes, side of cube 1 is a and side of cube 2 is b. and i have to find What is the ratio of: i)the area of their faces? ii)their volumes?